3. Oranges sold Price per Orange y (100s) x (cents) 6 10 4 20 5 30...

3.

Oranges sold

Price per Orange

y (100s)

x (cents)

6

10

4

20

5

30

4

40

3

50

1

60

Above are six weeks of data on the quantity of oranges sold (y) and price (x):

  1. Estimate the parameters of this equation.

  2. Predict the quantity of oranges sold if the price is 25 cents per orange.

  1. Derive the OLS estimators for 0 and 1 in the simple regression model.
Answer #1

For the given data on Oranges sold for each given price/Orange, set a regression equation in which the dependent variable is "Oranges sold y(100s)" and the independent variable is "Price per Orange x(cents)".

--------------------------------------------------------------------------------------------------------------------------------------------------------------

The result is:

SUMMARY OUTPUT Regression Statistics MultipleR R Square Adjusted R Square Standard Error Observations 0.899973247 0.809951846 0.762439807 0.839500986 ANOVA MS Significance F Regression Residual Total 1 12.01428571 12.01428571 17.0472973 0.014507625 4 2.8190476190.704761905 5 14.83333333 Coefficients Standard Error 6.733333333 0.781532885 8.615547036 0.0009976714.563450179 8.903216487 4.5634501798.903216487 0.0828571430.020067912-4.128837282 0.014507625 -0.138574599 -0.027139687 -0.138574599 -0.027139687 t Stat P-value Lower 95% Ue, 95% Lower 95.0% Uper 95.0% Intercept x (cents)

Thus, the intercept is 6.733333333 and the slope-coefficient is -0.082857143

It shall be noted that the intercept and the slope-coefficient are the parameters of the regression equation

y(100s) = 6.733333333 - 0.082857143 x

--------------------------------------------------------------------------------------------------------------------------------------------------------------

When x=25 cents

y(100s) = 6.733333333 - 0.082857143 * 25

=4.661904762

Hence, the quantity of Oranges sold is 4.661904762*100 = 466 Oranges

--------------------------------------------------------------------------------------------------------------------------------------------------------------

When price is 0 cent, the OLS estimator - intercept is 6.733333333 and expected y(100s) is 6.733333333

When price is 1 cent, the expected y(100s) is:

y(100s) = 6.733333333 - 0.082857143 * 1

=6.65047619